OpenStudy (anonymous):

a pair of shoes costing $62.50 is marked 40% off.What is the sale price of the shoes

OpenStudy (anonymous):

40/100 X 62.50 = 25 62.50-25 = 37.50

OpenStudy (amistre64):

P(1-%) = sale price 62.50(.60) = sale price

OpenStudy (amistre64):

10% = 6.25 ; times 6 is abt $37 and change

OpenStudy (anonymous):

37.50 62.50 - (40/100 X 62.50) = sale price